how do i rewrite the following trigonometric functions using angles between 0 and 90 degrees?
sin 150 degrees= sin________?
sin 150 degrees= cos________?
csc 45 degrees= sec________?
cos-30 degrees= cos________?
cos-30 degrees= sin________?
cot 75 degrees= tan________?

Answer by stanbon(57347)

how do i rewrite the following trigonometric functions using angles between 0 and 90 degrees?
sin 150 degrees= sin 30 degrees
----------------------------------
sin 150 degrees= cos 60 degrees
----------------------------------
csc 45 degrees= sec 45 degrees
----------------------------------
cos(-30) degrees= cos (30) degrees
----------------------------------
cos(-30) degrees= sin (60) degrees
------------------------------------
cot(75) degrees= tan (15) degrees
====================================
Cheers,
Stan H.